DX/x+y=dy/x+y=DX/-(x+y+2z)


1
Expert's answer
2021-05-10T17:10:48-0400

Given, "\dfrac{dx}{x+y}=\dfrac{dy}{x+y}=\dfrac{dx}{-(x+y+2z)}"


Taking first two terms-


"\dfrac{dx}{x+y}=\dfrac{dy}{x+y}"


"\Rightarrow dx=dy"


Integrate-

"x=y+c_1"


"c_1=x-y~~~-(1)"


Taking first and last term we get-


"\dfrac{dx}{x+y}=\dfrac{dx}{-(x+y+2z)}

\\[9pt]

\Rightarrow x+y=-x-y-2z

\\[9pt]

\Rightarrow z=-(x+y)"


Now Taking second and third term-


"\dfrac{dy}{x+y}=\dfrac{dz}{-(x+y+2(-x-y)}

\\[9pt]

\Rightarrow \dfrac{dy}{x+y}=\dfrac{dx}{x+y}

\\[9pt]

\Rightarrow dy=dx"

Integrate-


"y=x+c_2\\[9pt]

 c_2=y-x~~~~~-(2)"


The solution is-

"\phi(c_1,c_2)=0

\\[9pt]

\phi(x-y,y-x)=0"